Grade 12th passPhysical Chemistry

During the preparation of per sulfuric acid O2 gas is also released as by product when 9.72 litres of H2 releases at cathode and 2.35 litres of o2 at STP the weight of per disulphuric acid produced in grams is

Sumalika
8 Years agoGrade 12th pass

2 Answers

Arun
8 Years ago
Dear student
 
I am getting my answer as-
Weight of per disulphuric acid (H2S2O8) is 43.56 gm
Regards
Arun (askIITians forum expert)
Sanskar
4 Years ago
Solution 
Equivalent of H2,9208 + Equivalent of 02 (at anode) = Equivalent of 112 (at cathode)
Eq. of H2 = ( Volume released / Equivalent Volume)
                 = (9.72 /11.2)  =  0.87
 Eq. of 02 = ( 2.35 / 5.6)  =  0.42 
Wt. of H2S2O8 = g - eq x eq . wt = 0.45 x  (mol.wt / 2)
                                                     = (0.45 *194) / 2  
                                                     = 43.65